OpenStudy (anonymous):

write using exponents: a. 5sq.rt a^2 b. (3sq.rt b)^4 c. 4sq.rt 5t^3 d. 1/5sq. rt x^4

OpenStudy (anonymous):

a. \[\sqrt[5]{a ^{2}}\] b. \[(\sqrt[3]{b})^{4}\] c. \[\sqrt[4]{5t ^{3}}\] d. \[1/\sqrt[5]{x ^{4}}\]

OpenStudy (anonymous):

how did u get that?

OpenStudy (anonymous):

root means power 1/2........so wherever there is root substitute to power 1/2

OpenStudy (anonymous):

is that root of 5 and 3

OpenStudy (anonymous):

if it is for a answer is a^(2/5)